hii,,,i want to know how the relation

half life=0.693/(disintegration constant) is derived

Grade:12

2 Answers

Roshan Mohanty
10 years ago

For 1st order reactions
K=2.303 / t  * log(A/A-x)
when the reaction is half finished A-x=A/2
So K=2.303 / t(1/2)  * log(2)
      =0.693 / t(1/2) 

Vikas TU
10 years ago

 

PLZZZ    APPROVE!